Suppose a single electron orbits about a nucleus containing two protons (+2e), as would be the case for a helium atom from which one of the naturally occurring electrons is removed. The radius of the orbit is 4.68 × 10-11 m. Determine the magnitude of the electron's centripetal acceleration.
1
Expert's answer
2020-07-20T15:00:47-0400

The magnitude of the electron's centripetal acceleration can be defined from the Newton's second law:


"a = \dfrac{F}{m_e}"

where "F" is the force acting on the electron from the protons and "m_e = 9.1\times 10^{-31} kg" is the mass of electron.

The force "F" can be found using the Coulomb's law:


"F = k\dfrac{e\cdot 2p}{r^2}"

where "e =p= 1.6\times 10^{-19}C" is the charge of the electron and proton, "k = 9\times 10^{9} N\cdot m^2/C^2" is the Coulomb's constant and "r = 4.68\times 10^{-11}m" is the radius of the orbit.

Combining it together, obtain:


"a = k\dfrac{2pe}{m_er^2} = 9\times 10^9\dfrac{2\cdot 1.6^2\times 10^{-38}}{9.1\times 10^{-31}\cdot 4.68^2\cdot 10^{-22}} \approx 2.2\times 10^{23} m/s^2"

Answer. 2.2*10^23.